OpenStudy (anonymous):

Please help! I will give you a medal. :)) A student is running a 10-kilometer race. He runs 1 kilometer every 4 minutes. Select the function that describes his distance from the finish line. f(x) = - 1 over 4x + 4 f(x) = - 1 over 10x + 4 f(x) = - 1 over 10x + 10 f(x) = - 1 over 4x + 10

jimthompson5910 (jim_thompson5910):

if x is the time that has elapsed, then what should f(x) equal when x = 0?

OpenStudy (anonymous):

Um... 0?

jimthompson5910 (jim_thompson5910):

in other words, how far from the finish line should you be when the time is x = 0 seconds?

OpenStudy (anonymous):

Um... not distance from the finish line at all? :S

jimthompson5910 (jim_thompson5910):

hint: it's in the first sentence

jimthompson5910 (jim_thompson5910):

at the start of the race, you are _____ km from the finish line

OpenStudy (anonymous):

So he's 10 km from the finish line? So which function is that? :o

jimthompson5910 (jim_thompson5910):

because we know that f(x) = 10 when x = 0, we know that the y-intercept is 10

jimthompson5910 (jim_thompson5910):

so we go from y = mx+b to y = mx + 10

jimthompson5910 (jim_thompson5910):

how would you use the other given info to determine the slope? any ideas?

OpenStudy (anonymous):

I don't know. D:

jimthompson5910 (jim_thompson5910):

He runs 1 kilometer every 4 minutes so as x increases by 4, y decreases by 1 in other words rise = -1 run = 4 so the slope is ????

OpenStudy (anonymous):

umm... oh god this is stressful... umm

OpenStudy (anonymous):

lemme just try and do the math real quick i guess

jimthompson5910 (jim_thompson5910):

hint: slope = rise/run

OpenStudy (anonymous):

So... -0.25?? :o

jimthompson5910 (jim_thompson5910):

keep it as a fraction though

OpenStudy (anonymous):

what is -0.25 as a fraction??

OpenStudy (anonymous):

oh nvm. -1/4 ?? right?

jimthompson5910 (jim_thompson5910):

correct, so the function is \[\large f(x) = -\frac{1}{4}x + 10\]

OpenStudy (anonymous):

Ohh! Okay! The slope tip really helped~ Thanks! :DD

jimthompson5910 (jim_thompson5910):

you're welcome